# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

### v2

@@ III.3.2 Initial boundary formulation at a null hypersurface with $r(w,0)\neq 0$
If the areal distance function is non-trivial on the bounday $\mathcal{B}^{\epsilon}$, i.e. $r(w,0)\neq 0$, a different boundary value formulation will come about. Near the boundary the metric has the behaviour
@@
that casts (80) into a form similar to the expansion of (74). If that were the case, the initial boundary value problem could be discussed like in the previous section. It is indeed possible to find a coordinate transformation for (80) such that the resulting metric has $g_{AB}=\hat{\lambda}^{2}q_{AB}+O(\hat{\lambda}^{3})$, but such a transformation gives $g_{\hat{w}\hat{w}}|_{\hat{\lambda}=0}=O(1)\neq 0$ in the new coordinates. This violates the null condition at the boundary $\hat{\lambda}=0$ and cannot be removed without requiring $r_{(0),w}=0$. Since the coefficient $r_{(0)}$ corresponds to a mass, which becomes clear considering (73), we discard the restriction $r_{(0),w}=0$ as it implies that null boundary $\mathcal{B}^{\epsilon}$ has a mass which remains constant wrt a $w$ evolution. Thus, having $r_{(0)}\neq 0$ on $\mathcal{B}^{\epsilon}$ indicates that $\mathcal{B}^{\epsilon}$ is a null hypersurface with a given mass. The choice of (80) with $r_{(0)}>0$ corresponds to choosing a spherically symmetric Gaussian null coordinate system Moncrief and Isenberg (1983) at $\mathcal{B}^{\epsilon}$ since $g_{AB}$ is positive definite on $\mathcal{B}^{\epsilon}$. In the next section it will be shown that $\mathcal{B}^{\epsilon}$ with $r_{(0)}\neq 0$ can be related to a horizon of a black hole. Although there is no local coordinate transformation to cast (80) into a form like (74), it can still be simplified because $V_{(1)}$ can be set to zero everywhere on $\mathcal{B}^{\epsilon}$. The null vector generating $\mathcal{B}^{\epsilon}$ is $n^{a}\partial_{a}$ and the generating curves are non-affine null geodesics, which becomes clear when projecting (12) to $\mathcal{B}^{\epsilon}$ whose only component is
@@
If the coordinate transformation is chosen such that the term in the parenthesis of (83) vanishes, the non-affine null geodesics ruling $\mathcal{B}^{\epsilon}$ become affine null geodesics and $g_{\hat{w}\hat{w}}=O(\lambda^{2})$ near $\mathcal{B}^{\epsilon}$. This removal of the inaffinity is in fact a classical feature of non-affine geodesics Schouten (1954); Synge (1960); Poisson (2009) and not restricted to null geodesics. We make this choice in this section. In addition, we make the gauge choice $\alpha=0$ everwhere on $\mathcal{B}^{\epsilon}$. Evaluation of the supplementary equations (II) and (18) on $\mathcal{B}^{\epsilon}$ yields
@@
Equation (84c) can be used to determine $r_{,\lambda}$ everwhere on $\mathcal{B}^{\epsilon}$ provided $r$ and $Q$ are known on $\mathcal{B}^{\epsilon}$ and $r\neq 0$ on $\mathcal{B}^{\epsilon}$. The start values for solving the hypersurface equations (29d) and (29e) for $Z$ and $\mathcal{L}$ follow from evaluation of their definitions on $\mathcal{B}^{\epsilon}$
@@
which show that the initial values for $Z$ and $\mathcal{L}$ are known once $r$ and $\Phi_{,w}$ are known on $\mathcal{B}^{\epsilon}$. With the information collected above, a double null characteristic initial value problem can be formulated with a nontrivial conformal factor $r$ on the null boundary $\mathcal{B}^{\epsilon}$.
In a spherically symmetric spacetime, let $\mathscr{N}^{\epsilon}(w)$ be a family of null hypersurface $w=const$ with generators $\ell^{a}\partial_{a}$ that are affinely parameterized with an affine parameter $\lambda$. The affine parameter $\lambda$ is normalized with a parameter $\epsilon$ via $\ell^{a}\nabla_{a}\lambda=\epsilon$ with $\epsilon^{2}=1$. The binary parameter $\epsilon$ indicates whether a null hypersurface $w=const$ is an outgoing null hypersurface ($\epsilon=-1$) or an ingoing null hypersurface ($\epsilon=1$). The null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ intersect a null hypersurface $\mathcal{B}^{\epsilon}(w)$, where $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$. The generators of $\mathcal{B}^{\epsilon}(w)$ are null vectors $n^{a}\partial_{a}$. These generators obey the cross normalisation $\ell^{a}n_{a}=-1$ and the integral curves (null geodesics) of $n^{a}$ are affinely parameterized with the parameter $w$ on $\mathcal{B}^{\epsilon}(w)$. The intersections of $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ are spatial spherical cross sections $\Sigma(w)$ of co-dimension 2 with an intrinsic two metric $g_{AB}$. The volume element $\sqrt{\det(g_{AB})}=r^{2}(w,0)\sqrt{q}$ with $\sqrt{q}$ being the volume element of a unit sphere metric, $q_{AB}$, is nowhere zero on $\mathcal{B}^{\epsilon}(w)$ and for $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$. The function $r(w,0)$ is extended for values $\lambda>0$ on every null hypersurface $w=const$ to be be an areal distance so that the volume integral for a value $\lambda_{c}>0$ is given by $4\pi r^{2}(w,\lambda_{c})$. A solution of the Einstein-Maxwell-scalar field equations on the domain $\mathcal{B}^{\epsilon}(w)\cup\mathscr{N}^{\epsilon}(w)$ can be found for the double-null CIBVP with the following specifications:
  • provide a value for $\epsilon$ to specify the nature of the initial data surface $\mathscr{N}^{\epsilon}(0)$ with $w=0$
  • specify the value of the gauge charge $q$
  • on $\mathscr{N}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $F_{\Phi}(\lambda)$
  • on $\mathcal{B}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $\mathcal{N}_{\mathcal{B}}(w)$
@@
This (double null) CIBVP can be solved in the following manner
@@
    with the boundary data $\mathcal{N}_{\mathcal{B}}(w)$, the scalar field $\Phi(w,0)$ on $\mathcal{B}^{\epsilon}(w)$ and the initial values $r(0,0)$, $r_{,w}(0,0)$, $Q(0,0)$ and $r_{\lambda}(0,0)$ solve the hierarchical set of boundary equations
@@
    00 0 $=$ $Q_{,w}(w,0)+2i\pi qr^{2}(w,0)\left[\overline{\Phi}(w,0)\mathcal{N}_{\mathcal{B}}-\Phi(w,0)\overline{\mathcal{N}}_{\Phi}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;.$ (90)
@@
    $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda)$ (92)
@@
    on $\mathscr{N}^{\epsilon}(0)$, i.e. the null hypersurface $w=0$, integrate hierarchically the hypersurface equations (29a)-(29f) to obtain the fields $r(0,\lambda)$, $Q(0,\lambda)$, $\alpha(0,\lambda)$, $Z(0,\lambda)$, $\mathcal{L}(0,\lambda)$ and $V(0,\lambda)$ while using $\mathcal{F}_{\Phi}(\lambda)$, $\Phi(0,\lambda)$ and the respective boundary value fields evaluated at $w=0$.
@@
    with $r(0,\lambda)$, $\mathcal{L}(0,\lambda)$, $V(0,\lambda)$ and $\Phi_{,\lambda}(0,\lambda)$ algebraically calculate $\Phi_{,w}(0,\lambda)$ on $w=0$ using equation (29g), integrate this $w$ derivative (using for example a finite difference scheme) to find $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w>0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
@@
Comparing this formulation of the double null CIBVP with the one of the previous section Sec. III.3.1, reveals that the two are very similar. In fact, the only essential difference is step 2 - the determination of the nontrivial boundary fields in $\mathcal{B}^{\epsilon}(w)$. The reason for this difference is that for one of the formulations $r=0$ on $\mathcal{B}^{\epsilon}(w)$ while for the other one it is essentially non-trivial. The two different formulations give rise for different areas of application. The first scheme, where $r=0$ on $\mathcal{B}^{\epsilon}(w)$, is well suited for studying the evolution of initial data with a regular center. The second scheme is suited for the study of (charged) black holes in which $\mathcal{B}^{\epsilon}(w)$ is a horizon with an associated horizon mass of the black hole. Indeed, in Sec. V.1 the second scheme is used to derive the Reisner-Nordström solution of a charged black hole.
@@ IV Symmetry considerations
which is invariant under the action of the vector field $\xi$ that is the Lie derivative $\mathscr{L}_{\xi}z=\xi^{a}\partial_{a}z=0$. In principle we may choose $B=0$, because of the metric’s invariance of the coordinate transformation of the type (10). However, we will keep it general to allow a broader family of null affine coordinates. Inserting (95) into (94c) yields
@@
where $\hat{r}(\cdot)$ is any non-vanishing differentiable function. Finally, inserting (95) into (94d) gives us
@@
in which $\hat{V}(\cdot)$ is an arbitrary differentible function.
@@
The coordinate dependence of the metric potentials $r$ and $V$ for the different expressions for Killing vectors in Table 1 is general. For further study of particular applications it is practical to study the norm of the Killing vector either for large values of the affine parameter or its origin. The former will give us restrictions to the metric by requiring the Killing vector be asymptotically timelike, while the latter gives us necessary conditions, if the origin of the affine parameter is on a Killing horizon with a given surface gravity.
@@
$\xi^{a}\xi_{a}=-VA^{2}-\epsilon(A^{2})_{,w}\lambda+2\epsilon AB$ (101)
and in its asymptotic limit ($\lambda\rightarrow\infty$) we find using (36g), (37) and (LABEL:eq:Vinf)
@@
Hence, as seen in (IV), the asymptotic behavior of the Killing vector norm is determined by the asymptotic behavior of the areal distance through the functions $H$ and $R_{\infty}$. Therefore, in order to have a finite norm of this asymptotic limit, the linear term in $\lambda$ should vanish. One posibility is to set $A=0$, which from (101) implies a Killing vector field with vanishing norm of the form $\xi=B\partial_{\lambda}$. However, taking into account (98) we see that it implies that the vector field is zero everywhere or $\xi$ is a null Killing vector with $r=r(w)$ everywhere, which is inadmissible.
@@
If the Killing vector is required to be an asymptotically timelike unit Killing vector, one finds the condition
$-1=\left[2\epsilon B-\frac{c-2c\epsilon H(R_{\infty})_{,w}}{H}\right]cH$ (105)
@@
The relations (104) and (106) show that the components of a timelike Killing vector can be completely fixed by the asymptotic solution of the field equation. In particular, it is determined by the asymptotic behavior of the areal distance $r$ and an asymptotic time like Killing vector should have the form
$\begin{split}\xi=&cH(w)\partial_{w}+\\ &\left[-cH_{,w}\lambda+\frac{c\epsilon}{2H}-c(R_{\infty})_{,w}-\frac{\epsilon}{2cH}\right]\partial_{\lambda}+O(\lambda^{-1}).\end{split}$ (107)
In particular, regarding the choices of $A$ and $B$ in table 1 and requiring the Killing vectors to be asymptotically timelike unit Killing vectors implies using (98), (103) and (106).
@@
    … case $A=1$ and $B=0$, we find $c=1/H$. Moreover, since $r=r(\lambda)$, $H=const$ and $R_{\infty}=const$. Consequently, $\xi^{a}\xi_{a}=-1$ in the asymptotic limit can only be archived if $H=1$, i.e. the frame at large radii is a Bondi frame and the areal distance has the asymptotic expansion
@@
    …case $A=kw$ and $B=0$ gives $kw=cH(w)$. As Table 1 shows $r=r(w\lambda)$, the asymptotic expansion of $r$ in (36a) becomes $r(w\lambda)=\frac{kw\lambda}{c}+R_{\infty}(w)+O(\lambda^{-1})$ which requires $R_{\infty}=const$. But then the normalisation $\xi^{a}\xi_{a}=-1$ implies $c^{2}=1$, so that we can choose $c=1$. Consequently $r$ has the asymptotic expansion
@@
    …case $A=\frac{1}{2}w^{2}(w+k)^{-2}$ and $B=1$ gives ${w^{2}=2c(w+k)^{2}H}$. Then fixing $H$ in the asymptotic behavior of $r$ gives for (36a)
@@
    $R_{\infty}(w)=-cw+r_{\infty}\;\;,r_{\infty}=const.$ (111)
    To be consistient with the invariant, any solution for $r$ of the field equations with this Killing vector, which is in addition an asymptotical timelike unit vector, must have the large $\lambda$ expansion of the form
@@
    The determination of $c$ for the present case is postponed to V.1, because of the subtlety of the behavior of the Killing vector for $w=0$. There, it is a non-vanishing null tangent vector to the hypersurface, consequently, $\xi|_{w=0}=\partial_{\lambda}$. Requiring regularity of this surface (which implies regularity of $V$ given by (100) can be used to fix the value of the constant $c$ (see next section for an explicit example in an extremal Reissner-Nordströn black hole).
Of interest is also the behaviour of the norm of the Killing vector for $\lambda=0$, i.e on $\mathcal{B}^{\epsilon}$. At the origin ($\lambda=0$) of the affine parameter, the norm (101) takes the value
@@
If $V|_{\lambda=0}=0$ the surface $\lambda=0$ is a null hypersurface $\mathcal{B}^{\epsilon}$ and if $B=0$ and $A\neq 0$ the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon [^5] If $A=0$, we also have $\xi^{a}\xi_{a}=0$, but it is not a Killing horizon because only a vector field of the form $\xi\propto\partial_{w}$ can be a null generator of a Killing horizon at a null surface determined by $\lambda=0$. However, the possibility $A=B=0$ can be satisfied at an eventual bifurcating Killing horizon, where the Killing vector field vanishes.. Indeed, $B\neq 0$ in case iii) is crucial for the discussion of the (vanishing) surface gravity of an extremal black hole as is the case of extremal Reissner-Nordström black holes discussed in the next section. It turns out that for case iii) there also exists a Killing horizon not necessarily located at $\lambda=0$ but given implicitly by an expression of the form $F(w,\lambda)=0$. In the other cases, the corresponding surface gravity, $\kappa_{H}$, can be found from the invariant relation Wald (1984)

### v3

@@ III.3.2 Initial boundary formulation at a null hypersurface with $r(w,0)\neq 0$
If the areal distance function is nontrivial on the boundary $\mathcal{B}^{\epsilon}$, i.e., $r(w,0)\neq 0$, a different boundary value formulation will come about. Near the boundary the metric has the behavior
@@
that casts (80) into a form similar to the expansion of (74). If that were the case, the initial boundary value problem could be discussed as in the previous section. It is, indeed, possible to find a coordinate transformation for (80) such that the resulting metric has $g_{AB}=\hat{\lambda}^{2}q_{AB}+O(\hat{\lambda}^{3})$, but such a transformation gives $g_{\hat{w}\hat{w}}|_{\hat{\lambda}=0}=O(1)\neq 0$ in the new coordinates. This violates the null condition at the boundary $\hat{\lambda}=0$ and cannot be removed without requiring $r_{(0),w}=0$. Since the coefficient $r_{(0)}$ corresponds to a mass, which becomes clear considering (73), we discard the restriction $r_{(0),w}=0$ as it implies that null boundary $\mathcal{B}^{\epsilon}$ has a mass which remains constant with respect to a $w$ evolution. Thus, having $r_{(0)}\neq 0$ on $\mathcal{B}^{\epsilon}$ indicates that $\mathcal{B}^{\epsilon}$ is a null hypersurface with a given mass. The choice of (80) with $r_{(0)}>0$ corresponds to choosing a spherically symmetric Gaussian null coordinate system Moncrief and Isenberg (1983) at $\mathcal{B}^{\epsilon}$ since $g_{AB}$ is positive definite on $\mathcal{B}^{\epsilon}$. In the next section it will be shown that $\mathcal{B}^{\epsilon}$ with $r_{(0)}\neq 0$ can be related to a horizon of a black hole. Although there is no local coordinate transformation to cast (80) into a form like (74), it can still be simplified because $V_{(1)}$ can be set to zero everywhere on $\mathcal{B}^{\epsilon}$. The null vector generating $\mathcal{B}^{\epsilon}$ is $n^{a}\partial_{a}$ and the generating curves are nonaffine-null geodesics, which becomes clear when projecting (12) to $\mathcal{B}^{\epsilon}$ whose only component is
@@
If the coordinate transformation is chosen such that the term in the parentheses of (83) vanishes, the nonaffine-null geodesics ruling $\mathcal{B}^{\epsilon}$ become affine-null geodesics and $g_{\hat{w}\hat{w}}=O(\lambda^{2})$ near $\mathcal{B}^{\epsilon}$. This removal of the inaffinity is, in fact, a classical feature of nonaffine geodesics Schouten (1954); Synge (1960); Poisson (2009) and not restricted to null geodesics. We make this choice in this section. In addition, we make the gauge choice $\alpha=0$ everywhere on $\mathcal{B}^{\epsilon}$. Evaluation of the supplementary equations (II) and (18) on $\mathcal{B}^{\epsilon}$ yields
@@
Equation (84c) can be used to determine $r_{,\lambda}$ everywhere on $\mathcal{B}^{\epsilon}$ provided $r$ and $Q$ are known on $\mathcal{B}^{\epsilon}$ and $r\neq 0$ on $\mathcal{B}^{\epsilon}$. The initial values for solving the hypersurface equations (29d) and (29e) for $Z$ and $\mathcal{L}$ follow from their definitions on $\mathcal{B}^{\epsilon}$
@@
which show that the initial values for $Z$ and $\mathcal{L}$ are known once $r$ and $\Phi_{,w}$ are known on $\mathcal{B}^{\epsilon}$. With the information collected above, a double-null CIBVP can be formulated with a nontrivial conformal factor $r$ on the null boundary $\mathcal{B}^{\epsilon}$.
In a spherically symmetric spacetime, let $\mathscr{N}^{\epsilon}(w)$ be a family of null hypersurface $w=$const with generators $\ell^{a}\partial_{a}$ that are affinely parametrized with an affine parameter $\lambda$. The affine parameter $\lambda$ is normalized with a parameter $\epsilon$ via $\ell^{a}\nabla_{a}\lambda=\epsilon$ with $\epsilon^{2}=1$. The binary parameter $\epsilon$ indicates whether a null hypersurface $w=$const is an outgoing null hypersurface ($\epsilon=-1$) or an ingoing null hypersurface ($\epsilon=1$). The null hypersurfaces $\mathscr{N}^{\epsilon}(w)$ intersect a null hypersurface $\mathcal{B}^{\epsilon}(w)$, where $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$. The generators of $\mathcal{B}^{\epsilon}(w)$ are null vectors $n^{a}\partial_{a}$. These generators obey the cross normalization $\ell^{a}n_{a}=-1$ and the integral curves (null geodesics) of $n^{a}$ are affinely parametrized with the parameter $w$ on $\mathcal{B}^{\epsilon}(w)$. The intersections of $\mathscr{N}^{\epsilon}(0)$ and $\mathcal{B}^{\epsilon}(w)$ are spatial spherical cross sections $\Sigma(w)$ of codimension-2 with an intrinsic two metric $g_{AB}$. The volume element $\sqrt{\det(g_{AB})}=r^{2}(w,0)\sqrt{q}$ with $\sqrt{q}$ being the volume element of a unit sphere metric, $q_{AB}$, is nowhere zero on $\mathcal{B}^{\epsilon}(w)$ and for $\lambda=0$ on $\mathscr{N}^{\epsilon}(w)$. The function $r(w,0)$ is extended for values $\lambda>0$ on every null hypersurface $w=$const to be an areal distance so that the volume integral for a value $\lambda_{c}>0$ is given by $4\pi r^{2}(w,\lambda_{c})$. A solution of the Einstein-Maxwell-scalar field equations on the domain $\mathcal{B}^{\epsilon}(w)\cup\mathscr{N}^{\epsilon}(w)$ can be found for the double-null CIBVP with the following specifications:
  • provide a value for $\epsilon$ to specify the nature of the initial data surface $\mathscr{N}^{\epsilon}(0)$ with $w=0$;
  • specify the value of the gauge charge $q$;
  • on $\mathscr{N}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $F_{\Phi}(\lambda)$;
  • on $\mathcal{B}^{\epsilon}(0)$ specify a free (sufficiently differentiable) complex function $\mathcal{N}_{\mathcal{B}}(w)$;
@@
This double-null CIBVP can be solved in the following manner
@@
    With the boundary data $\mathcal{N}_{\mathcal{B}}(w)$, the scalar field $\Phi(w,0)$ on $\mathcal{B}^{\epsilon}(w)$, and the initial values $r(0,0)$, $r_{,w}(0,0)$, $Q(0,0)$ and $r_{\lambda}(0,0)$ solve the hierarchical set of boundary equations
@@
    00 0 $=$ $Q_{,w}(w,0)+2i\pi qr^{2}(w,0)\left[\overline{\Phi}(w,0)\mathcal{N}_{\mathcal{B}}-\Phi(w,0)\overline{\mathcal{N}}_{\Phi}\right]\bigg{|}_{\mathcal{B}^{\epsilon}}\;\;,$ (90)
@@
    $\Phi_{,\lambda}(0,\lambda)=\mathcal{F}_{\Phi}(\lambda),$ (92)
@@
    on $\mathscr{N}^{\epsilon}(0)$, i.e. the null hypersurface $w=0$, integrate hierarchically the hypersurface equations (29a)-(29f) to obtain the fields $r(0,\lambda)$, $Q(0,\lambda)$, $\alpha(0,\lambda)$, $Z(0,\lambda)$, $\mathcal{L}(0,\lambda)$, and $V(0,\lambda)$, while using $\mathcal{F}_{\Phi}(\lambda)$, $\Phi(0,\lambda)$ and the respective boundary value fields evaluated at $w=0$.
@@
    With $r(0,\lambda)$, $\mathcal{L}(0,\lambda)$, $V(0,\lambda)$, and $\Phi_{,\lambda}(0,\lambda)$ algebraically calculate $\Phi_{,w}(0,\lambda)$ on $w=0$ using Eq. (29g), integrate this $w$ derivative (using, for example, a finite difference scheme) to find $\Phi(\Delta w,\lambda)$ at some time $w=\Delta w\neq 0$ on a null hypersurface $\mathscr{N}^{\epsilon}(\Delta w)$
@@
Comparing this formulation of the double-null CIBVP with the one of the previous section Sec. III.3.1, reveals that the two are very similar. In fact, the only essential difference is step 2 – the determination of the nontrivial boundary fields in $\mathcal{B}^{\epsilon}(w)$. The reason for this difference is that for one of the formulations $r=0$ on $\mathcal{B}^{\epsilon}(w)$, while for the other one it is essentially nontrivial.
The two different formulations give rise for different areas of application. The first scheme, where $r=0$ on $\mathcal{B}^{\epsilon}(w)$, is well suited for studying the evolution of initial data with a regular center. The second scheme is suited for the study of (charged) black holes in which $\mathcal{B}^{\epsilon}(w)$ is a horizon with an associated horizon mass of the black hole. Indeed, in Sec. V.1 the second scheme is used to derive the Reissner-Nordström solution of a charged black hole.
@@ IV Symmetry considerations
which is invariant under the action of the vector field $\xi$ that is the Lie derivative $\mathscr{L}_{\xi}z=\xi^{a}\partial_{a}z=0$. In principle, we may choose $B=0$, because of the metric’s invariance of the coordinate transformation of the type (10). However, we will keep it general to allow a broader family of affine-null coordinates. Inserting (95) into (94c) yields
@@
where $\hat{r}(\cdot)$ is any nonvanishing differentiable function. Finally, inserting (95) into (94d) gives us
@@
in which $\hat{V}(\cdot)$ is an arbitrary differentiable function.
@@
The coordinate dependence of the metric potentials $r$ and $V$ for the different expressions for Killing vectors in Table 1 is general. For further particular applications, it is practical to study the norm of the Killing vector either for large values of the affine parameter or at its origin. The former will give us restrictions to the metric by requiring the Killing vector be asymptotically timelike, while the latter gives us necessary conditions, if the origin of the affine parameter is on a Killing horizon with a given surface gravity.
@@
$\xi^{a}\xi_{a}=-VA^{2}-\epsilon(A^{2})_{,w}\lambda+2\epsilon AB.$ (101)
In its asymptotic limit ($\lambda\rightarrow\infty$), we find using (36g), (37) and (38):
@@
Hence, as seen in (IV), the asymptotic behavior of the Killing vector norm is determined by the asymptotic behavior of the areal distance through the functions $H$ and $R_{\infty}$. Therefore, in order to have a finite norm of this asymptotic limit, the linear term in $\lambda$ should vanish. One possibility is to set $A=0$, which from (101) implies the Killing vector field $\xi=B\partial_{\lambda}$ with vanishing norm. However, taking into account (98), we see that it implies that the vector field is zero everywhere or $\xi$ is a null Killing vector with $r=r(w)$ everywhere, which is inadmissible.
@@
If the Killing vector is required to be an asymptotically timelike unit vector, one finds the condition
$-1=\left[2\epsilon B-\frac{c-2c\epsilon H(R_{\infty})_{,w}}{H}\right]cH,$ (105)
@@
The relations (104) and (106) show that the components of a timelike Killing vector can be completely fixed by the asymptotic solution of the field equation. In particular, it is determined by the asymptotic behavior of the areal distance $r$. An asymptotic timelike Killing vector should therefore have the form
$\begin{split}\xi=&cH(w)\partial_{w}\\ &+\left[-cH_{,w}\lambda+\frac{c\epsilon}{2H}-c(R_{\infty})_{,w}-\frac{\epsilon}{2cH}\right]\partial_{\lambda}+O(\lambda^{-1}).\end{split}$ (107)
In particular, regarding the choices of $A$ and $B$ in Table 1 and requiring the Killing vectors to be asymptotically timelike unit vectors implies using (98), (103) and (106):
@@
    For case $A=1$ and $B=0$, we find $c=1/H$. Moreover, since $r=r(\lambda)$, $H$=const and $R_{\infty}$=const. Consequently, $\xi^{a}\xi_{a}=-1$ in the asymptotic limit can only be achieved if $H=1$; i.e. the frame at large radii is a Bondi frame and the areal distance has the asymptotic expansion
@@
    The case $A=kw$ and $B=0$ gives $kw=cH(w)$. As Table 1 shows $r=r(w\lambda)$, the asymptotic expansion of $r$ in (36a) becomes $r(w\lambda)=\frac{kw\lambda}{c}+R_{\infty}(w)+O(\lambda^{-1})$, which requires $R_{\infty}$=const. But then the normalization $\xi^{a}\xi_{a}=-1$ implies $c^{2}=1$, so that we can choose $c=1$. Consequently $r$ has the asymptotic expansion
@@
    The case $A=\frac{1}{2}w^{2}(w+k)^{-2}$ and $B=1$ gives ${w^{2}=2c(w+k)^{2}H}$. Then fixing $H$ in the asymptotic behavior of $r$ gives for (36a)
@@
    $R_{\infty}(w)=-cw+r_{\infty}\;\;,r_{\infty}=\mathrm{const}.$ (111)
    To be consistent with the invariant, any solution for $r$ of the field equations with this Killing vector, which is in addition an asymptotical timelike unit vector, must have the large $\lambda$ expansion of the form
@@
    The determination of $c$ for the present case is postponed to Sec. V.1, because of the subtlety of the behavior of the Killing vector for $w=0$. There, it is a nonvanishing null tangent vector to the hypersurface, consequently, $\xi|_{w=0}=\partial_{\lambda}$. A regularity requirement of this surface [which implies regularity of $V$ given by (100)] can be used to fix the value of the constant $c$ (see next section for an explicit example in an extremal Reissner-Nordström black hole).
Of interest is also the behavior of the norm of the Killing vector for $\lambda=0$, i.e on $\mathcal{B}^{\epsilon}$. At the origin ($\lambda=0$) of the affine parameter, the norm (101) takes the value
@@
If $V|_{\lambda=0}=0$ the surface $\lambda=0$ is a null hypersurface $\mathcal{B}^{\epsilon}$, and if $B=0$ and $A\neq 0$ the null hypersurface $\mathcal{B}^{\epsilon}$ is a Killing horizon [^8] If $A=0$, we also have $\xi^{a}\xi_{a}=0$, but it is not a Killing horizon because only a vector field of the form $\xi\propto\partial_{w}$ can be a null generator of a Killing horizon at a null surface determined by $\lambda=0$. However, the possibility $A=B=0$ can be satisfied at an eventual bifurcating Killing horizon, where the Killing vector field vanishes.. Indeed, $B\neq 0$ in case (iii) is crucial for finding the (vanishing) surface gravity of an extremal black hole as is the case of extremal Reissner-Nordström black holes discussed in the next section. It turns out that for case iii) there also exists a Killing horizon not necessarily located at $\lambda=0$ but given implicitly by an expression of the form $F(w,\lambda)=0$. In the other cases, the corresponding surface gravity, $\kappa_{H}$, can be found from the invariant relation Wald (1984)